# Fuzzy Classification Aggregation

### v1

@@ 3 Results
Despite its technical flavor, Theorem 3 has as a particular case the class of Weighted Arithmetic Means, and also any other power mean with an odd exponent. [^7] A power mean with an odd exponent is a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ such that $f(x_{1},\ldots,x_{n})=(\frac{1}{n}\sum_{i=1}^{n}x_{i}^{m})^{\frac{1}{m}}$ where $m$ is an odd number. Proposition $1$ of Cailloux et al., (2024) is a particular case of Theorem 3, where condition 1 is implied by the complementary condition of their result.
A more general result, valid for $m\geq p\geq 2$ can be attained if we slightly change the surjectivity conditions. For this general case, we consider fuzzy classifications $c:X\rightarrow\mathbb{R}$, and the surjectivity requirements are such that for all $x\in X$, and for a given $s\in\mathbb{R}$, it is the case that $\sum_{t=1}^{p}c(x)_{t}=s$, and that $\sum_{j=1}^{m}c(x_{j})_{t}\geq s$ if $s\geq 0$ for all $t\in P$, or $\sum_{j=1}^{m}c(x_{j})_{t}\leq s$ if $s<0$ for all $t\in P$. We can think of this setting as $n$ individuals assigning $s$ hours of use of $m$ machines into $p$ different tasks that need $s$ hours to be finished, or $s$ euros that $m$ persons have to spend in $p$ different projects that need at least $s$ euros to be done. We use $\mathcal{C}^{\ast}$ to denote this set of classifications. Thus, the ${}^{\ast}$ is a mapping $\alpha:{\mathcal{C}^{\ast}}^{n}\rightarrow\mathcal{C}^{\ast}$ such that $\alpha(\bm{c})(x_{j})$ indicates, for example, the different hours assigned to the different tasks of the machine $x_{j}$.
It is easy to see that when $m=p$, all the inequalities of the surjectivity conditions turn into equalities. This implies that trivially, the ${}^{\ast}$ satisfies the $k$-allocation property for any $k=m\geq 2$ (Aczél and Wagner,, 1980), that is, if $\sum_{j=1}^{m}c(x_{j})_{t}=s$, then $(\sum_{j=1}^{m}(\alpha(\bm{c})(x_{j}))_{1},\ldots,\sum_{j=1}^{m}(\alpha(\bm{c})(x_{j}))_{p})=\bm{s}$ where $\bm{s}$ is a $p-$dimensional vector with an $s$ on every entry. The $k$-allocation property allows us to characterize the symmetric ${}^{\ast}$s that satisfy Fuzzy Consensus for any value of $m$ such that $m\geq p\geq 2$.
@@
Let $m\geq p\geq 2$. A symmetric ${}^{\ast}$ satisfies Fuzzy Consensus if, and only if, it is a Weighted Arithmetic Mean.
@@
Aczél and Wagner, (1980) show that a mapping $\bm{c}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ satisfies the k-allocation property for $k=2,3$ and is bounded if, and only if, it is a Weighted Arithmetic Mean. In our setting, symmetry implies that for a given category $t\in P$, we have the same elementary ${}^{\ast}$ for every object. By Fuzzy Consensus, they are bounded and thus they are Weighted Arithmetic Means. By a similar proof to the one used in Theorem 2, it is easy to see that the elementary ${}^{\ast}$s are the same for every category, concluding our proof. ∎
@@ 4 Final Remarks
In this work, we present an analysis from a fuzzy point of view, of the challenge of classifying $m$ objects into $p$ different categories. The classifications of the objects by the agents and the rules are no longer crisp statements about to which category the object is assigned. Instead of that, the classifications are expressed in terms of degrees of assignment to each of the categories. In the crisp setting for more than two objects and two categories, requiring the rule to fill each category with at least one object, to be unanimous and independent, implies the existence of an agent such that objects are classified according to the opinions of that agent (Maniquet and Mongin,, 2016). Even with a weaker surjectivity condition, this result can not have a major improvement (Cailloux et al.,, 2024, showing the existence of an essential dictator).
The fuzzy setting proves advantageous, as the surjectivity conditions can assume diverse yet natural interpretations. Strong interpretations, akin to those in Alcantud et al., (2019), can extend the scope of impossibility results. However, by considering weaker surjectivity conditions and different versions of Unanimity, we can circumvent these limitations. Our findings show that rules that belong to the family of Weighted Arithmetic Means are the only ones that satisfy Zero Unanimity and Independence. Under different interpretations of the classification process and the consensual axioms, we can obtain characterization results that hold for $m\geq p\geq 2$.

### v3

@@ 3 Results
Despite its technical flavour, Proposition 1 has as a particular case the class of Weighted Arithmetic Means, and also any other power mean with an odd exponent. [^8] A power mean with an odd exponent is a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ such that $f(x_{1},\ldots,x_{n})=(\frac{1}{n}\sum_{i=1}^{n}x_{i}^{m})^{\frac{1}{m}}$ where $m$ is an odd number. Proposition $1$ of Cailloux et al., (2024) is a particular case of Proposition 1, where condition 1 is implied by the complementary condition of their result.
A more general result, valid for $m\geq p\geq 2$ can be attained if we slightly change the surjectivity conditions. For this general case, we consider fuzzy classifications $c^{\ast}:X\rightarrow\mathbb{R}^{p}$, and the surjectivity requirements are such that for all $x\in X$, and for a given $s\in\mathbb{R}$, it is the case that $\sum_{t=1}^{p}c^{\ast}(x)_{t}=s$, and that $\sum_{j=1}^{m}c^{\ast}(x_{j})_{t}\geq s$ if $s\geq 0$ for all $t\in P$, or $\sum_{j=1}^{m}c^{\ast}(x_{j})_{t}<s$ if $s<0$ for all $t\in P$. We can think of this setting as $n$ operators assigning $s$ hours of use of $m$ machines into $p$ different tasks that need $s$ hours to be finished, or $s$ euros that $m$ persons have to spend in $p$ different projects that need at least $s$ euros to be done. We use $\mathcal{C}^{\ast}$ to denote this set of classifications. Thus, the FCAF is a mapping $\alpha:{\mathcal{C}^{\ast}}^{N}\rightarrow\mathcal{C}^{\ast}$ such that $\alpha(\bm{c}^{\ast})(x_{j})$ indicates, for example, the different hours assigned to the different tasks of the machine $x_{j}$.
We have that an optimal FCAF satisfies the $k$-allocation property for any $k=m\geq 2$ (Aczél and Wagner,, 1980), that is, if for any $t\in P$ such that $\sum_{j=1}^{m}c^{\ast}(x_{j})_{t}=s$ for all $i\in N$, it is the case that $\sum_{j=1}^{m}\alpha(\bm{c}^{\ast})(x_{j})_{t}=s$. The $k$-allocation property allows us to characterize the optimal and symmetric FCAFs that satisfy Fuzzy Consensus for any value of $m$ such that $m\geq p\geq 2$.
@@
Let $m\geq p\geq 2$. An optimal and symmetric FCAF satisfies Fuzzy Consensus and Non-Dictatorship if, and only if, it is a Weighted Arithmetic Mean with a non-degenerate set of weights.
@@
Aczél and Wagner, (1980) show that a mapping $\bm{c}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ satisfies the k-allocation property for $k=2,3$ and is bounded if, and only if, it is a Weighted Arithmetic Mean. In our setting, symmetry implies that for a given category $t\in P$, we have the same elementary FCAF for every object. By optimality, the k-allocation property is satisfied by the elementary FCAFs. Fuzzy Consensus implies they are bounded and thus they are Weighted Arithmetic Means, and finally, Non-Dictatorship implies that the set of weights is non-degenerate. By a similar proof to the one used in Theorem 1, it is easy to see that the elementary FCAFs are the same for every category, concluding our proof. ∎
@@ 4 Final Remarks
In this work, we present an analysis from a fuzzy point of view, of the challenge of classifying $m$ objects into $p$ different categories. The classifications of the objects by the agents and the rules are no longer crisp statements about to which category the object is assigned. Instead of that, the classifications are expressed in terms of degrees of assignment to each of the categories. In the crisp setting for more than two objects and two categories, requiring the rule to fill each category with at least one object, to be independent and unanimous, implies the existence of an agent such that objects are classified according to the opinions of that agent (Maniquet and Mongin,, 2016). Even with a weaker unanimity condition, this result can not have a major improvement (Cailloux et al.,, 2024, showing the existence of an essential dictator).
The fuzzy setting proves advantageous, as the surjectivity conditions can assume diverse yet natural interpretations. Strong interpretations, akin to those in Alcantud et al., (2019), can extend the scope of impossibility results. However, by considering weaker surjectivity conditions and different versions of Unanimity, we can circumvent these limitations. Our findings show that rules that belong to the family of Weighted Arithmetic Means are the only ones that satisfy Optimality, Independence, and Zero Unanimity. Under different interpretations of the classification process and the consensual axioms, we can obtain characterization results that hold for $m\geq p\geq 2$.